OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (nirmalnema):

3x − 6y = 3 multiply by 7 7x − 5y = −11 mutiply by 3 new equations. 21x-42y=21 21x-15y=-33 21+42y=15y-33 7+14y=5y-11 14y-5y=-7-11 9y=-18 y=-2